Serene Sound produces a high quality audio tape used in the recording industry.  Serene allocates variable overhead to production at a rate of $12 per batch manufactured.  The company’s monthly fixed overhead costs average $72,000.  An average of 500 batches per month is considered normal.  During June, Serene produced 450 batches of audio tape and incurred actual overhead costs of $79,500.

1          Compute the following amounts:

a          Total overhead applied to production in June amounted to $__________.

b          Total overhead budgeted in June for the 450 batches manufactured amounted to $__________.

c          Serene’s overhead spending variance was $__________ (favorable/unfavorable).

d          Serene’s overhead volume variance was $__________ (favorable/unfavorable).
Answer #1

Answer to:-

Fixed Overhead applied rate = total budgeted fixed overhead/ Budgeted quantity.

   72000/500 = $144

Variable overhead applied rate = $ 12 (Given in question)

a) Budgeted rate * Actual quantity  

(72000/500*450)+ (12*450) = $70200

b) Actual quantity * variable overhead rate + Standard fixed overheads

(450*12)+72000 = $ 77400

c ) Actual overheads - (Actual quantity * budgeted rate

79500-(450*((72000/500)+12) = $ 9300 Favorable

d) (Actual quantity * budgeted rate) - (budgeted quantity * budgeted rate

(450*156) -(500*156) $ 7800 Unfavorable